gvgtftyfk:

can someone help me What value of x is a solution to x2+2=102

mxddi3:

So we have \[x^2+2=102\] First we want to get the variable on a side by itself. So, we are going to subtract 2 from both sides. This leaves us with: \[x^2=100\] Now, since it's x^2, we are going to take the SQUARE ROOT of both sides.

NotLeiland:

102-2 = 100 sqrroot of 100 = answer